# Anosov representations, strongly convex cocompact groups and weak eigenvalue gaps

### v3

@@ 1.3. Characterizations of strongly convex cocompact groups
Anosov representations of hyperbolic groups are closely related with real projective geometry and geometric structures. Fix an integer $d\geqslant 3$. A subset $\Omega$ of the projective space $\mathbb{P}(\mathbb{R}^{d})$ is called properly convex if it is contained in an affine chart on which $\Omega$ is bounded and convex. The domain $\Omega$ is called strictly convex if it is properly convex and $\partial\Omega$ does not contain projective line segments.
Let $\Gamma$ be a discrete subgroup of $\mathsf{PGL}_{d}(\mathbb{R})$ which preserves a properly convex domain $\Omega$ of $\mathbb{P}(\mathbb{R}^{d})$. The full orbital limit set $\Lambda_{\Omega}(\Gamma)$ of $\Gamma$ in $\Omega$ is the set of accumulation points of all $\Gamma$-orbits in $\partial\Omega$ (see [17, Def. 1.10]). The group $\Gamma$ acts convex cocompactly on $\Omega$ if the convex hull of $\Lambda_{\Omega}(\Gamma)$ in $\Omega$ is non-empty and has compact quotient by $\Gamma$ (see [17, Def. 1.11]). The group $\Gamma$ is called strongly convex cocompact in $\mathbb{P}(\mathbb{R}^{d})$ if it acts convex cocompactly on some properly convex domain $\Omega$ with strictly convex and $C^{1}$-boundary. The work of Danciger–Guéritaud–Kassel [17] and independently of Zimmer [39], shows that Anosov representations can be essentially (up to composition with a Lie group homomorphism) viewed as convex cocompact actions on properly convex domains in some real projective space. We refer the reader to [17, Thm. 1.4 & 1.15] and [39, Thm. 1.22 & 1.25]. There are also related results in the setting of naively convex cocompact groups, see [24, Thm. 1.13].
@@
([17, Thm. 1.4]) Let $\Gamma$ be an infinite discrete subgroup of $\mathsf{PGL}_{d}(\mathbb{R})$ which preserves a properly convex domain of $\mathbb{P}(\mathbb{R}^{d})$. Then $\Gamma$ is strongly convex cocompact in $\mathbb{P}(\mathbb{R}^{d})$ if and only if $\Gamma$ is word hyperbolic and the natural inclusion $\Gamma\xhookrightarrow{}\mathsf{PGL}_{d}(\mathbb{R})$ is $P_{1}$-Anosov.
@@
(i) $\Gamma$ is strongly convex cocompact in $\mathbb{P}(\mathbb{R}^{d})$.
(ii) The inclusion $\Gamma\xhookrightarrow{}\mathsf{PGL}_{d}(\mathbb{R})$ is a quasi-isometric embedding, $\Gamma$ preserves a properly convex domain $\Omega$ of $\mathbb{P}(\mathbb{R}^{d})$ and there exists a $\Gamma$-invariant closed convex subset $\mathcal{C}$ of $\Omega$ such that $\big{(}\mathcal{C},d_{\Omega}\big{)}$ is Gromov hyperbolic.
The previous theorem generalizes the well known fact that a discrete subgroup $\Gamma$ of $\mathsf{PO}(d,1)$, $d\geq 2$, is convex cocompact if and only if the inclusion $\Gamma\xhookrightarrow{}\mathsf{PO}(d,1)$ is a quasi-isometric embedding.
@@
We also intoduce a definition of a Gromov product on $G\times G$ which we use for the proof of Theorem 1.10 (see Lemma 8.1). Let us remark that there are similar definitions of Gromov products in [6, §3] and [7, §8], however these are defined on the flag spaces associated to $G$. The Gromov product from [7] is also vector valued into a Cartan subspace of the Lie algebra of $G$.
@@
$\big{(}g\cdot h)_{\varphi}:=\frac{1}{4}\big{\langle}\varphi,\mu(g)+\mu(g^{-1})+\mu(h)+\mu(h^{-1})-\mu(g^{-1}h)-\mu(h^{-1}g)\big{\rangle}.$
@@
Let $G$ be a real semisimple Lie group, fix $\theta\subset\Delta$ a subset of simple restricted roots of $G$ and let $\{\omega_{\alpha}\}_{\alpha\in\theta}$ be the associated set of fundamental weights. Suppose that $\Gamma$ is a word hyperbolic group and $\rho:\Gamma\rightarrow G$ is a $P_{\theta}$-Anosov representation. There exist $C,c>1$ with the property that for every $\alpha\in\theta$ and $\gamma,\delta\in\Gamma$ we have
$C^{-1}(\gamma\cdot\delta)_{e}-c\leqslant\big{(}\rho(\gamma)\cdot\rho(\delta)\big{)}_{\omega_{\alpha}}\leqslant C(\gamma\cdot\delta)_{e}+c.$
@@
We prove Proposition 1.12 as follows: by [22, Prop. 1.8] any semisimplification $\rho^{ss}$ of $\rho$ is $P_{\theta}$-Anosov and hence, by using Lemma 2.11, we may replace $\rho$ with $\rho^{ss}$. Then we compare the Gromov product relative to the fundamental weight $\{\omega_{\alpha}\}_{\alpha\in\theta}$ with the Gromov product with respect to the Hilbert metric $d_{\Omega}$ for some properly convex domain and then use Theorem 1.9.
@@
We first explain how Theorem 1.1 is related to the equivalence $(3)\Leftrightarrow(5)$ in [29, Thm. 1.7], see also [27, Thm. 5.47]. A subgroup $\Gamma$ of a real reductive Lie group $G$ is called $\tau_{\textup{mod}}$-asymptotically embedded [29, Def. 6.12], if it is $\tau_{\textup{mod}}$-regular, $\tau_{\textup{mod}}$-antipodal, word hyperbolic and there exists a $\Gamma$-equivariant homeomorphism $\nu:\partial_{\infty}\Gamma\rightarrow\Lambda_{\tau_{\textup{mod}}}(\Gamma)$. Here $\tau_{\textup{mod}}$ corresponds to the choice of a subset of simple restricted roots $\eta\subset\Delta$ of $G$, $\tau_{\textup{mod}}$-antipodal means that the map $\nu$ is transverse to itself i.e. for $x\neq y$ the pair $(\nu(x),\nu(y))$ is transverse and $\tau_{\textup{mod}}$-regular corresponds to $P_{\eta}$-divergence.
Theorem 1.1 follows from a theorem of Kapovich–Leeb–Porti [29, Thm. 1.7] in the case where both limit maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}$ satisfy the Cartan property (see Definition 4.1). Under this assumption, there exists $\rho$-equivariant embedding $\xi:\partial_{\infty}\Gamma\rightarrow G/P$ with $P=P_{\theta}^{+}\cap P_{\theta^{\star}}^{+}$, where ${}^{\star}:\Delta\rightarrow\Delta$ denotes the opposition involution and $\theta^{\star}=\{\alpha^{\star}:\alpha\in\theta\}$. Note that the pair of maps $(\xi^{+},\xi^{-})$ is compatible and transverse, hence $\xi$ is injective. The map $\xi$ satisfies the Cartan property, maps onto the $\tau_{\textup{mod}}$-limit set $\Lambda_{\tau_{\textup{mod}}}(\rho(\Gamma))$ hence $\rho(\Gamma)$ is $\tau_{\textup{mod}}$-asymptotically embedded and the assumptions of [29, Thm. 1.7] are satisfied.
We also remark that Guichard-Guéritaud-Kassel-Wienhard proved in [22, Thm. 1.3, (1$\Leftrightarrow$(2)] that a representation $\rho:\Gamma\rightarrow G$ is $P_{\theta}$-Anosov if and only if $\rho$ is $P_{\theta}$-divergent and admits a pair of continuous, $\rho$-equivariant, dynamics preserving and transverse maps $\xi^{\pm}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{\pm}$. Theorem 1.1 follows by [22, Thm. 1.3, (1$\Leftrightarrow$(2)] under the additional assumption that both limit maps are dynamics preserving.
@@
In §2 we provide the necessary background from Lie theory, hyperbolic groups and the Floyd boundary and recall Labourie’s dynamical definition of Anosov representations. In §3 we prove some preliminary results which we use for the proof of Theorem 1.1. In §4 we define the Cartan property for an equivariant map $\xi:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{\pm}$ and discuss the uniform gap summation property of [22] in the more general setting of finitely generated groups. In §5 we discuss (strong) property (U) and prove Theorem 1.6 and Corollary 1.8. In §6 we define a Gromov product for a representation $\rho$ and prove that is comparable with the usual Gromov product on the domain group when $\rho$ is Anosov. Next, in §7 we prove Theorem 1.1 and in §8 we give the proof of Theorem 1.10. In §9 we provide conditions for the direct product of two representations to be Anosov. Finally, in §10 we provide examples of discrete and faithful representations of surface groups showing that the assumptions of our main results are necessary.
@@ Acknowledgements.
I would like to thank Richard Canary and Fanny Kassel for their support during the course of this work and for their comments in previous versions of this paper. I would also like to thank Michael Kapovich, Giuseppe Martone, Andrés Sambarino and Feng Zhu for helpful discussions. The author was partially supported by grants DMS-1564362 and DMS-1906441 from the National Science Foundation as well as from the European Research Council (ERC) under the European’s Union Horizon 2020 research and innovation programme (ERC starting grant DiGGeS, grant agreement No 715982).
@@ 2. Background
In this section, we recall definitions from Lie theory, review several facts for hyperbolic groups, the Floyd boundary, provide Labourie’s dynamical definition of Anosov representations and also discuss several facts for semisimple representations. We mainly follow the notation from [22, §2].
@@
Throughout this paper $\Gamma$ is a finitely generated group equipped with a finite generating subset $S$, inducing a left invariant word metric $d_{\Gamma}$ on the Cayley graph $C_{\Gamma}$ of $\Gamma$. For $\gamma\in\Gamma$ we set $|\gamma|_{\Gamma}:=d_{\Gamma}(\gamma,e)$. A linear representation $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$, $d\geq 2$, is called irreducible if $\rho(\Gamma)$ does not preserve any proper subspace of $\mathbb{R}^{d}$. The representation $\rho$ is called strongly irreducible if for every finite-index subgroup $H$ of $\Gamma$ the restriction $\rho|_{H}$ is irreducible.
@@ 2.1. Lie theory
We will always consider $G$ to be a semisimple Lie subgroup of $\mathsf{SL}_{m}(\mathbb{R})$, $m\geqslant 2$, of non-compact type with finitely many connected components. The Zariski topology on $G$ is the subspace topology induced from real algebraic subsets of $\mathsf{SL}_{m}(\mathbb{R})$.
We fix a maximal compact subgroup $K$ of $G$, unique up to conjugation, a Cartan decomposition $\mathfrak{g}=\mathfrak{t}\oplus\mathfrak{p}$ where $\mathfrak{t}=\textup{Lie}(K)$, $\mathfrak{p}$ is the orthogonal complement of $\mathfrak{t}$ with respect to the Killing form on $\mathfrak{g}$, and the Cartan subspace $\mathfrak{a}$ which is a maximal abelian subalgebra of $\mathfrak{g}$ contained in $\mathfrak{p}$. The real rank of $G$ is the dimension of $\mathfrak{a}$ as a real vector space.
For a linear form $\beta\in\mathfrak{a}^{\ast}$ we shall use the notation $\left\langle\beta,H\right\rangle=\beta(H)$ for $H\in\mathfrak{a}$. There is a decomposition of $\mathfrak{g}$ into the common eigenspaces of the transformations $X\mapsto[H,X],H\in\mathfrak{a}$, called the restricted root decomposition
@@
where $\mathfrak{g}_{\alpha}=\big{\{}X\in\mathfrak{g}:[H,X]=\langle\alpha,H\rangle X\ \forall H\in\mathfrak{a}\big{\}}$ and $\Sigma=\big{\{}\alpha\in\mathfrak{a}^{\ast}:\mathfrak{g}_{\alpha}\neq 0\big{\}}$ is the set of restricted roots of $G$. We fix $H_{0}\in\mathfrak{a}$ with $\alpha(H_{0})\neq 0$ for every $\alpha\in\Sigma$. Denote by $\Sigma^{+}=\big{\{}\alpha\in\Sigma:\langle\alpha,H_{0}\rangle>0\big{\}}$ the set of positive roots and fix $\Delta\subset\Sigma^{+}$ the simple positive roots. For any simple restricted root $\alpha\in\Delta$, denote by $\omega_{\alpha}$ the fundamental weight with respect to $\alpha\in\Delta$, e.g. see [22, §3.1].
@@
and denote by $P_{\theta}^{\pm}=N_{G}(\mathfrak{p}_{\theta}^{\pm})$. A subgroup $P$ of $G$ is parabolic if it normalizes some parabolic subalgebra. Two parabolic subgroups $P^{+}$ and $P^{-}$ of $G$ are called opposite if there $\theta\subset\Delta$ and $g\in G$ such that $P^{\pm}=gP_{\theta}^{\pm}g^{-1}$.
Let $\overline{\mathfrak{a}}^{+}:=\big{\{}H\in\mathfrak{a}:\langle\alpha,H\rangle\geqslant 0,\ \forall\alpha\in\Delta\big{\}}$. There exists a decomposition
@@
and $\mu(g)$ denotes the Cartan projection of $g$. The map $\mu:G\rightarrow\overline{\mathfrak{a}}^{+}$ is continuous and proper and is called the Cartan projection. The Lyapunov projection $\lambda:G\rightarrow\overline{\mathfrak{a}}^{+}$ is the map defined as follows for $g\in G$,
@@
An element $g\in G$ is called $P_{\theta}$-proximal if $\min_{\alpha\in\theta}\left\langle\alpha,\lambda(\rho(\gamma))\right\rangle>0$. Equivalently, $g$ has two fixed points $x_{g}^{+}\in G/P_{\theta}^{+}$ and $V_{g}^{-}\in G/P_{\theta}^{-}$ such that the pair $(x_{g}^{+},V_{g}^{-})$ is transverse and for every $x\in G/P_{\theta}^{+}$ transverse to $V_{g}^{-}$, we have $\lim_{n}g^{n}x=x_{g}^{+}$. The element $g$ is called $P_{\theta}$-biproximal if $g$ and $g^{-1}$ are both $P_{\theta}$-proximal and we denote by $x_{g}^{-}$ the attracting fixed point of $g^{-1}$ in $G/P_{\theta}^{-}$.
@@
The case of $G=\mathsf{SL}_{d}(\mathbb{R})$. We denote by $(e_{1},\ldots,e_{d})$ the canonical basis of $\mathbb{R}^{d}$ and set $e_{j}^{\perp}:=\oplus_{j\neq i}\mathbb{R}e_{j}$. The group $\mathsf{SO}(d)=\{g\in\mathsf{SL}_{d}(\mathbb{R}):gg^{t}=I_{d}\big{\}}$ is the unique, up to conjugation, maximal compact subgroup of $\mathsf{SL}_{d}(\mathbb{R})$. A Cartan subspace for $\mathfrak{g}$ is the subspace $\mathfrak{a}=\textup{diag}_{0}(d)$ of all diagonal matrices with zero trace. Let $\varepsilon_{i}\in\mathfrak{a}^{\ast}$ be the projection to the $(i,i)$-entry. The closed dominant Weyl chamber of $\mathfrak{a}$ is $\overline{\mathfrak{a}}^{+}:=\big{\{}\textup{diag}(a_{1},\ldots,a_{d}):a_{1}\geqslant\ldots\geqslant a_{d},\ \sum_{i=1}^{d}a_{i}=0\big{\}}$ and we have the Cartan decomposition $\mathsf{SL}_{d}(\mathbb{R})=\mathsf{SO}(d)\exp(\overline{\mathfrak{a}}^{+})\mathsf{SO}(d)$. The restricted root decomposition is $\mathfrak{sl}_{d}(\mathbb{R})=\mathfrak{a}\oplus\bigoplus_{i\neq j}\mathbb{R}E_{ij}$, where $E_{ij}$ denotes the $d\times d$ elementary matrix with $1$ at the $(i,j)$ entry and 00 everywhere else. The set of restricted roots is $\big{\{}\varepsilon_{i}-\varepsilon_{j}:i\neq j\big{\}}$ and of simple positive roots $\big{\{}\varepsilon_{i}-\varepsilon_{i+1}:i=1,\ldots,d-1\big{\}}$. For each $i=1,\ldots,d-1$, the associated fundamental weight is $\omega_{\varepsilon_{i}-\varepsilon_{i+1}}=\sum_{k=1}^{i}\varepsilon_{k}$. For an element $g\in\mathsf{SL}_{d}(\mathbb{R})$ we denote by $\sigma_{i}(g)$ and $\ell_{i}(g)$ the $i$-th singular value and eigenvalue of $g$. Recall the connection between eigenvalues and singular values $\sigma_{i}(g)=\sqrt[]{\ell_{i}(gg^{t})}$. The Cartan and Lyapunov projections of $g\in\mathsf{SL}_{d}(\mathbb{R})$ are

### v4

@@ 1.3. Characterizations of strongly convex cocompact groups
Anosov representations of hyperbolic groups are closely related to real projective geometry and geometric structures. Fix an integer $d\geqslant 3$. A subset $\Omega$ of the projective space $\mathbb{P}(\mathbb{R}^{d})$ is called properly convex if it is contained in an affine chart on which $\Omega$ is bounded and convex. The domain $\Omega$ is called strictly convex if it is properly convex and $\partial\Omega$ does not contain projective line segments.
Let $\Gamma$ be a discrete subgroup of $\mathsf{PGL}_{d}(\mathbb{R})$ which preserves a properly convex domain $\Omega$ of $\mathbb{P}(\mathbb{R}^{d})$. The full orbital limit set $\Lambda_{\Omega}(\Gamma)$ of $\Gamma$ in $\Omega$ is the set of accumulation points of all $\Gamma$-orbits in $\partial\Omega$ (see [19, Def. 1.10]). The group $\Gamma$ acts convex cocompactly on $\Omega$ if the convex hull of $\Lambda_{\Omega}(\Gamma)$ in $\Omega$ is non-empty and has compact quotient by $\Gamma$ (see [19, Def. 1.11]). The group $\Gamma$ is called strongly convex cocompact in $\mathbb{P}(\mathbb{R}^{d})$ if it acts convex cocompactly on some properly convex domain $\Omega$ with strictly convex and $C^{1}$-boundary. The work of Danciger–Guéritaud–Kassel [19] and independently of Zimmer [44], shows that Anosov representations can be essentially (up to composition with a Lie group homomorphism) viewed as convex cocompact actions on properly convex domains in some real projective space. We refer the reader to [19, Thm. 1.4 & 1.15] and [44, Thm. 1.22 & 1.25]. There are also related results in the more broad setting of naively convex cocompact groups, see [26, Thm. 1.13].
@@
([19, Thm. 1.4]) Let $\Gamma$ be an infinite discrete subgroup of $\mathsf{PGL}_{d}(\mathbb{R})$ which preserves a properly convex domain of $\mathbb{P}(\mathbb{R}^{d})$. Then $\Gamma$ is strongly convex cocompact in $\mathbb{P}(\mathbb{R}^{d})$ if and only if $\Gamma$ is word hyperbolic and the natural inclusion $\Gamma\xhookrightarrow{}\mathsf{PGL}_{d}(\mathbb{R})$ is $P_{1}$-Anosov.
@@
  1. (i)
    $\Gamma$ is strongly convex cocompact in $\mathbb{P}(\mathbb{R}^{d})$.
  2. (ii)
    The inclusion $\Gamma\xhookrightarrow{}\mathsf{PGL}_{d}(\mathbb{R})$ is a quasi-isometric embedding, $\Gamma$ preserves a properly convex domain $\Omega$ of $\mathbb{P}(\mathbb{R}^{d})$ and there exists a $\Gamma$-invariant closed convex subset $\mathcal{C}$ of $\Omega$ such that $\big(\mathcal{C},d_{\Omega}\big)$ is Gromov hyperbolic.
The previous theorem generalizes the well-known fact that a discrete subgroup $\Gamma$ of $\mathsf{PO}(d,1)$, $d\geq 2$, is convex cocompact if and only if $\Gamma\xhookrightarrow{}\mathsf{PO}(d,1)$ is a quasi-isometric embedding.
@@
We also intoduce a definition of a Gromov product on $G\times G$ which we use for the proof of Theorem 1.10 (see Lemma 8.1). Let us remark that there are similar notions of Gromov products in [8, §3] and [9, §8] defined on appropriate flag spaces of $G$. The Gromov product from [9] is also vector valued into a Cartan subspace of the Lie algebra of $G$.
@@
$\big(g\cdot h)_{\varphi}:=\frac{1}{4}\varphi\Big(\mu(g)+\mu(g^{-1})+\mu(h)+\mu(h^{-1})-\mu(g^{-1}h)-\mu(h^{-1}g)\Big).$
@@
Let $G$ be a real semisimple Lie group, fix $\theta\subset\Delta$ a subset of simple restricted roots of $G$ and let $\{\omega_{\alpha}\}_{\alpha\in\theta}$ be the associated set of fundamental weights. Suppose that $\Gamma$ is a word hyperbolic group and $\rho:\Gamma\rightarrow G$ is a $P_{\theta}$-Anosov representation. There exist $C,c>1$ with the property that for every $\alpha\in\theta$ and $\gamma_{1},\gamma_{2}\in\Gamma$ we have
$C^{-1}(\gamma_{1}\cdot\gamma_{2})_{e}-c\leqslant\big(\rho(\gamma_{1})\cdot\rho(\gamma_{2})\big)_{\omega_{\alpha}}\leqslant C(\gamma_{1}\cdot\gamma_{2})_{e}+c.$
@@
We prove Proposition 1.12 as follows: by [24, Prop. 1.8] any semisimplification $\rho^{ss}$ of $\rho$ is $P_{\theta}$-Anosov and hence, by using Lemma 2.11, we may replace $\rho$ with $\rho^{ss}$. Then we compare the Gromov product relative to the fundamental weight $\{\omega_{\alpha}\}_{\alpha\in\theta}$ with the Gromov product with respect to the Hilbert metric $d_{\Omega}$ for some properly convex domain and then use Theorem 1.9.
@@
We first explain how Theorem 1.1 is related to the equivalence $(3)\Leftrightarrow(5)$ in [31, Thm. 1.7], see also [29, Thm. 5.47]. A subgroup $\Gamma$ of a real reductive Lie group $G$ is called $\tau_{\textup{mod}}$-asymptotically embedded [31, Def. 6.12], if it is $\tau_{\textup{mod}}$-regular, $\tau_{\textup{mod}}$-antipodal, word hyperbolic and there exists a $\Gamma$-equivariant homeomorphism $\nu:\partial_{\infty}\Gamma\rightarrow\Lambda_{\tau_{\textup{mod}}}(\Gamma)$. Here $\tau_{\textup{mod}}$ corresponds to the choice of a subset of simple restricted roots $\eta\subset\Delta$ of $G$, $\tau_{\textup{mod}}$-antipodal means that the map $\nu$ is transverse to itself i.e. for $x\neq y$ the pair $(\nu(x),\nu(y))$ is transverse and $\tau_{\textup{mod}}$-regular corresponds to $P_{\eta}$-divergence.
Theorem 1.1 follows from a theorem of Kapovich–Leeb–Porti [31, Thm. 1.7] in the case where both limit maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}$ satisfy the Cartan property (see Definition 4.1). Under this assumption, there exists $\rho$-equivariant embedding $\xi:\partial_{\infty}\Gamma\rightarrow G/P$ with $P=P_{\theta}^{+}\cap P_{\theta^{\star}}^{+}$, where $^{\star}:\Delta\rightarrow\Delta$ denotes the opposition involution and $\theta^{\star}=\{\alpha^{\star}:\alpha\in\theta\}$. Note that the pair of maps $(\xi^{+},\xi^{-})$ is compatible and transverse, hence $\xi$ is injective. The map $\xi$ satisfies the Cartan property, maps onto the $\tau_{\textup{mod}}$-limit set $\Lambda_{\tau_{\textup{mod}}}(\rho(\Gamma))$ hence $\rho(\Gamma)$ is $\tau_{\textup{mod}}$-asymptotically embedded and the assumptions of [31, Thm. 1.7] are satisfied.
[…]
@@
In §2 we provide the necessary background from Lie theory, hyperbolic groups and the notion of the Floyd boundary and recall Labourie’s dynamical definition of Anosov representations. In §3 we prove some preliminary results which we use for the proof of Theorem 1.1. In §4 we define the Cartan property for an equivariant map $\xi:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{\pm}$ and discuss the uniform gap summation property of [24] in the more general setting of finitely generated groups. In §5 we discuss (strong) property (U) and prove Theorem 1.6 and Corollary 1.8. In §6 we define a Gromov product for a representation $\rho$ and prove that is comparable with the usual Gromov product on the domain group when $\rho$ is Anosov. Next, in §7 we prove Theorem 1.1 and in §8 we give the proof of Theorem 1.10. In §9 we provide conditions for the direct product of two representations to be Anosov. Finally, in §10 we provide examples of discrete and faithful representations of surface groups showing that the assumptions of our main results are necessary.
@@ Acknowledgements.
I would like to thank Richard Canary and Fanny Kassel for their support during the course of this work and for their comments in previous versions of this paper. I would also like to thank Sami Douba, Michael Kapovich, Giuseppe Martone, Andrés Sambarino and Feng Zhu for helpful discussions. Finally, I would also like to thank the anonymous referee for carefully reading the paper and their comments and suggestions that improved the exposition. The author was partially supported by grants DMS-1564362 and DMS-1906441 from the National Science Foundation as well as from the European Research Council (ERC) under the European’s Union Horizon 2020 research and innovation programme (ERC starting grant DiGGeS, grant agreement No 715982).
@@ 2. Background
In this section, we recall definitions from Lie theory, review several facts for hyperbolic groups, the Floyd boundary, provide Labourie’s dynamical definition of Anosov representations and also discuss several facts for semisimple representations. We mainly follow the notation from [24, §2].
@@
Throughout this paper $\Gamma$ is a finitely generated group equipped with a finite generating subset $S$, inducing a left invariant word metric $d_{\Gamma}$ on the Cayley graph $C_{\Gamma}$ of $\Gamma$. For $\gamma\in\Gamma$ we set $|\gamma|_{\Gamma}:=d_{\Gamma}(\gamma,e)$, where $e\in\Gamma$ is the identity element. A linear representation $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$, $d\geq 2$, is called irreducible if $\rho(\Gamma)$ does not preserve any non-trivial proper vector subspace of $\mathbb{R}^{d}$. The representation $\rho$ is called strongly irreducible if for every finite-index subgroup $H$ of $\Gamma$ the restriction $\rho|_{H}$ is irreducible. We equip the vector space $\mathbb{R}^{d}$ with the canonical basis $(e_{1},\ldots,e_{d})$, where $e_{i}$ is the vector with $1$ on the $i^{\textup{th}}$ coordiante and zero everywhere else, and the standard Euclidean inner product $\langle\cdot,\cdot\rangle$. For a subspace $V\subset\mathbb{R}^{d}$, $V^{\perp}=\{v\in\mathbb{R}^{d}:\langle v,v^{\prime}\rangle=0,\forall v^{\prime}\in V\}$ is the orthogonal complement of $V$.
@@ 2.1. Lie theory
We will always consider $G$ to be a semisimple Lie subgroup of $\mathsf{SL}_{m}(\mathbb{R})$, $m\geqslant 2$, of non-compact type with finitely many connected components. The Zariski topology on $G$ is the subspace topology induced from the Zariski topology on $\mathsf{SL}_{m}(\mathbb{R})$.
We fix a maximal compact subgroup $K$ of $G$, unique up to conjugation, a Cartan decomposition $\mathfrak{g}=\mathfrak{t}\oplus\mathfrak{p}$ where $\mathfrak{t}=\textup{Lie}(K)$, $\mathfrak{p}$ is the orthogonal complement of $\mathfrak{t}$ with respect to the Killing form on $\mathfrak{g}$, and the Cartan subspace $\mathfrak{a}\subset\mathfrak{g}$ which is a maximal abelian subalgebra of $\mathfrak{g}$ contained in $\mathfrak{p}$. The real rank of $G$ is the dimension of $\mathfrak{a}$ as a real vector space.
There is a decomposition of $\mathfrak{g}$ into the common eigenspaces of the transformations $X\mapsto[H,X],H\in\mathfrak{a}$, called the restricted root decomposition
@@
where $\mathfrak{g}_{\alpha}=\big\{X\in\mathfrak{g}:[H,X]=\alpha(H)X,\forall H\in\mathfrak{a}\big\}$ and $\Sigma=\big\{\alpha\in\mathfrak{a}^{\ast}:\mathfrak{g}_{\alpha}\neq 0\big\}$ is the set of restricted roots of $G$. Fix $H_{0}\in\mathfrak{a}$ with $\alpha(H_{0})\neq 0$ for every $\alpha\in\Sigma$. Denote by $\Sigma^{+}=\big\{\alpha\in\Sigma:\alpha(H_{0})>0\big\}$ the set of positive roots and fix $\Delta\subset\Sigma^{+}$ the simple positive roots. For any simple restricted root $\alpha\in\Delta$, denote by $\omega_{\alpha}$ the fundamental weight with respect to $\alpha\in\Delta$, see [24, §3.1].
@@
and denote by $P_{\theta}^{\pm}=N_{G}(\mathfrak{p}_{\theta}^{\pm})$. A subgroup $P$ of $G$ is parabolic if it normalizes some parabolic subalgebra. A pair of parabolic subgroups $(P^{+},P^{-})$ of $G$ are called opposite if there $\theta\subset\Delta$ and $g\in G$ such that $(P^{+},P^{-})=(gP_{\theta}^{+}g^{-1},gP_{\theta}^{-}g^{-1})$.
Let $\overline{\mathfrak{a}}^{+}:=\big\{H\in\mathfrak{a}:\alpha(H)\geqslant 0,\forall\alpha\in\Delta\big\}$. There exists a decomposition
@@
and $\mu(g)\in\overline{\mathfrak{a}}^{+}$ denotes the Cartan projection of $g$. The map $\mu:G\rightarrow\overline{\mathfrak{a}}^{+}$ is called the Cartan projection and is continuous and proper. The Lyapunov projection $\lambda:G\rightarrow\overline{\mathfrak{a}}^{+}$ is the map defined as follows for $g\in G$,
@@
An element $g\in G$ is called $P_{\theta}$-proximal if $\min_{\alpha\in\theta}\alpha(\lambda(g)))>0$. Equivalently, $g$ has two fixed points $x_{g}^{+}\in G/P_{\theta}^{+}$ and $V_{g}^{-}\in G/P_{\theta}^{-}$ such that the pair $(x_{g}^{+},V_{g}^{-})$ is transverse and for every $x\in G/P_{\theta}^{+}$ transverse to $V_{g}^{-}$, we have $\lim_{n}g^{n}x=x_{g}^{+}$. The element $g$ is called $P_{\theta}$-biproximal if $g$ and $g^{-1}$ are both $P_{\theta}$-proximal and we denote by $x_{g}^{-}$ the attracting fixed point of $g^{-1}$ in $G/P_{\theta}^{-}$.
For a matrix $h=(h_{ij})_{i,j=1}^{d}$ in $\mathsf{GL}_{d}(\mathbb{R})$ its transpose is $h^{t}:=(h_{ji})_{i,j=1}^{d}$.
@@
The case of $G=\mathsf{SL}_{d}(\mathbb{R})$. Recall that $(e_{1},\ldots,e_{d})$ denotes the canonical basis of $\mathbb{R}^{d}$ and $e_{j}^{\perp}:=\bigoplus_{j\neq i}\mathbb{R}e_{j}$. The group $\mathsf{SO}(d)=\{g\in\mathsf{SL}_{d}(\mathbb{R}):gg^{t}=I_{d}\big\}$ is the unique, up to conjugation, maximal compact subgroup of $\mathsf{SL}_{d}(\mathbb{R})$. A Cartan subspace for $\mathfrak{g}$ is the subspace $\mathfrak{a}=\textup{diag}_{0}(d)$ of all diagonal matrices with zero trace. Let $\varepsilon_{i}\in\mathfrak{a}^{\ast}$ be the projection to the $(i,i)$-entry. The closed dominant Weyl chamber of $\mathfrak{a}$ is $\overline{\mathfrak{a}}^{+}:=\big\{\textup{diag}(a_{1},\ldots,a_{d}):a_{1}\geqslant\ldots\geqslant a_{d},\ \sum_{i=1}^{d}a_{i}=0\big\}$ and we have the Cartan decomposition $\mathsf{SL}_{d}(\mathbb{R})=\mathsf{SO}(d)\exp(\overline{\mathfrak{a}}^{+})\mathsf{SO}(d)$. The restricted root decomposition is $\mathfrak{sl}_{d}(\mathbb{R})=\mathfrak{a}\oplus\bigoplus_{i\neq j}\mathbb{R}E_{ij}$, where $E_{ij}$ denotes the $d\times d$ elementary matrix with $1$ at the $(i,j)$ entry and 0 everywhere else. The set of restricted roots is $\big\{\varepsilon_{i}-\varepsilon_{j}:i\neq j\big\}$ and of simple positive roots $\big\{\varepsilon_{i}-\varepsilon_{i+1}:i=1,\ldots,d-1\big\}$. For each $i=1,\ldots,d-1$, the associated fundamental weight is $\omega_{\varepsilon_{i}-\varepsilon_{i+1}}=\sum_{k=1}^{i}\varepsilon_{k}$. For an element $g\in\mathsf{SL}_{d}(\mathbb{R})$ we denote by $\sigma_{i}(g)$ and $\ell_{i}(g)$ the $i$-th singular value and modulus of eigenvalues of $g$. Recall the connection between moduli of eigenvalues and singular values $\sigma_{i}(g)=\sqrt[]{\ell_{i}(gg^{t})}$. The Cartan and Lyapunov projections of $g\in\mathsf{SL}_{d}(\mathbb{R})$ respectively are